OpenStudy (anonymous):

find any real numbers for which the rational expression is undefined. 6x^3+7 ________ 4x^2+24x

OpenStudy (anonymous):

the denominator is never zero, so it is defined for all real numbers

OpenStudy (marissalovescats):

Are you trying to simplify 6x^3-7/4x^2+24x??

OpenStudy (anonymous):

notice that for all \(x\) you have \(x^2\geq 0\) and so \(4x^2+24\geq 24\) meaning the denominator cannot be zero, so you have no worries here about dividing by zero

OpenStudy (anonymous):

what would the answer be? im confued

OpenStudy (anonymous):

Satellite, I think you didn't see the x in 24x in the denominator. Isn't the denominator zero when x =0 or -6? \[\frac{ 6x^3 -7 }{ 4x^2 + 24x } = \frac{ 6x^3 -7 }{ 4x(x+6) }\]